OpenStudy (anonymous):

A low energy electromagnetic wave would also have A. low frequency and long wavelength B.high frequency and long wavelength C.high frequency and short wavelength D.low frequency and short wavelength PLEASE HELP FAN YOU AND MEDAL

OpenStudy (irishboy123):

\(E = h \nu = \frac{h \ c}{\lambda}\)

OpenStudy (anonymous):

IS D

OpenStudy (irishboy123):

or A or B or C :p \(Energy = h \times frequency = \frac{h \times speed \ of \ light}{wavelength}\) if E is "low", what would you need in terms of frequency and in terms of wavelength?

OpenStudy (irishboy123):

mmm :p http://openstudy.com/study#/updates/557f3eece4b07028ea605b81

OpenStudy (theeric):

Haha, this is also a more physics related question. I think @IrishBoy123 is right, both with his theory and the answer being A or B or C or D.